OpenStudy (anonymous):

an airplane is flying relative to the air at 100m/s in the direction east. the wind velocity relative to the ground is 30m/s at an angle of 45 degrees north of east. what is the speed of the plane relative to the ground?

OpenStudy (anonymous):

\[V_{b}=V _{a}+V _{b/a}\]

OpenStudy (anonymous):

Do you know how to make Velocity Triangle?

OpenStudy (anonymous):

1. Vpa = 100m/s. @ 0 deg. Vag =30m/s @ 45 deg,CCW. Vpg = Vpa + Vag, Vpg = (100 + 30cos45) + i30sin45, Vpg = 121 + i21.2, Vpg=sqrt((121)^2 + (21.2)^2)=122.9 m/s.